OpenStudy (anonymous):

Multiply and simplify: √6(√3+5√2)

OpenStudy (psymon):

Well, we can use this property to help us solve it: \[\sqrt{a}*\sqrt{b}=\sqrt{ab} \]So when you distribute the sqrt6 in, you get this: \[\sqrt{6}*\sqrt{3}+5\sqrt{2}*\sqrt{6} \]BEcause of our square root properties, this can become: \[\sqrt{18}+5\sqrt{12}\]From here the best we can do is factor those square roots. we want to factor them in such a way that they will reduce. In order to do this, we look for factors that are perfect squares. So doing that we find: \[\sqrt{18} = \sqrt{2}*\sqrt{9} = 3\sqrt{2}\]And as for the 5 sqrt12 part: \[5\sqrt{3}*\sqrt{4}=10\sqrt{3}\]So all together, this simplifies down to: \[3\sqrt{2} + 10\sqrt{3}\]

OpenStudy (anonymous):

thank u =)